# Dwarf Galaxies in the MATLAS Survey: Hubble Space Telescope Observations of Nuclear Star Clusters

### v1

@@ 3 HST Observations
In this study, a subsample of 79 of the 2210 dwarf galaxies identified in the MATLAS survey were observed with the HST Advanced Camera for Surveys (ACS) in the Wide Field Channel (WFC) mode. The HST observations were performed during the Cycle 27 program GO-16082 (PI: O. Müller), as well as the Cycle 28 and 29 snapshot programs SNAP-16257 and SNAP-16711 (PI: F. Marleau). We use the reduced CTE-corrected (Charge Transfer Efficiency corrected) drc drizzled images produced by the STScI standard pipeline, that have a pixel size of 0.05 arcsec/pixel. Each dwarf is centered in one of the charge coupled devices (CCDs) to optimize the background estimate and ensure that no object is lost (e.g., GCs) in the chip gap or outside the field of view. The galaxies were observed in the F606W and F814W filters, with an exposure time in each filter of 1030 s for MATLAS-2019, and 824 s for the others. We computed the VEGAmag zero-points from the python module acszpt222https://acstools.readthedocs.io/en/latest/acszpt.html made by the ACS Team. The targets of the follow-up observations were selected to study the GC population of the MATLAS UDGs (e.g., Müller et al. 2020, 2021; Marleau et al. 2021, 2024; Pfeffer et al. 2024; Buzzo et al. 2025; Haacke et al. 2025). Thus, our sample is mainly composed of low surface brightness and large dwarf galaxies (see Figure 2 from Marleau et al. 2024).
@@ 4 The nucleated dwarf sample
In Poulain et al. (2021), we defined a nucleus as a compact source located up to 0.5 effective radius from the photocenter of the dwarf, and that is the brightest compact source within one effective radius. Based on this definition, and after removing one likely foreground star contaminant using the Gaia DR2 catalog (Gaia Collaboration et al. 2018), 507 nucleated dwarfs were identified in the MATLAS dwarf galaxy sample. We note, however, that of the nucleus sample are located within a galactocentric distance of 0.1 effective radius, or 0.83″, a value lower than the observed PSF of the ground based observation of the MATLAS data. For 16 nucleated dwarfs, up to three compact sources of similar magnitude were located close to the photocenter and were brighter than the surrounding point sources. In these cases, we defined those galaxies as multinucleated.
Of the 507 identified MATLAS nucleated dwarfs, 30 have HST observations. However, the ACS camera having a significantly higher resolution than the MegaCam camera, we are now able to marginally resolve the nuclei, leading to the rejection of two nuclei that were actually background sources, and the identification of 15 new nuclei. Two of these are newly resolved double-nuclei located in already flagged nucleated galaxies. This results in a sample of 41 nucleated dwarfs with HST data to be studied. In Figure 1, we show a comparison between the resolution of the central region from the HST and MATLAS images for six nucleated, including three of these new nuclei, and color cutouts of the 41 nucleated dwarfs observed with HST are displayed in Figure 2. In terms of morphology, 40 of the nucleated dwarfs were classified as dE during the initial classification (Habas et al. 2020), and only one MATLAS-682 was defined as a dI.
@@ 5.1 Sérsic modeling
In order to obtain the structural parameters and photometry of the NSCs, we divided the modeling process into two parts. The aim of the first step was to remove the smooth light profile of the dwarf galaxy from the image. As the background is not constant all over the HST image, we estimated the local background of the galaxy. We computed the median sky from clipping in an annulus of radius 15″ on the outskirts of the galaxy – where the light profile stops decreasing to become constant. To optimize the galaxy modeling, we masked all nuclei with a circle of radius 1″, a good compromise to cover each NSC without masking a too large region of the galaxies. If necessary, we manually created an additional mask of surrounding bright sources, such as galaxies and stars, with the help of saoimageds9 (Smithsonian Astrophysical Observatory 2000). Finally, we ran galfit (Peng et al. 2010) to apply a Sérsic profile (Sérsic 1963) to the dwarf galaxy, fixing the value of the sky to the estimate. The Sérsic model has long been used in the literature for the NSC modeling (e.g., Graham & Spitler 2009; Seth et al. 2010; Turner et al. 2012; Lyubenova et al. 2013; Carson et al. 2015; Spengler et al. 2017). Because of the exposure time of the observations, it can be very difficult to differentiate the observed galaxies from the background given their low-surface brightness. As a consequence, the Sérsic models systematically have a smaller Re than the models on the CFHT images (Poulain et al. 2021), as the flux in the outer part of the galaxies blends into the sky noise. We also obtained less accurate photometric measurements, especially for the faintest galaxies. As a consequence, when the galaxies structural properties and photometry are studied, we used the values from Poulain et al. (2021) based on the CFHT images. The second step focused on the Sérsic modeling of the NSCs. For this, we created a cutout of size 5″ centered on the nucleus from the galaxy cleaned image, that is, the residual image from galfit. The PSF varies with the position on the CCD and the stellar population of the stars, thus, to precisely compute the structural properties of the NSCs, we generated the PSF of each NSC from tinytim (Krist et al. 2011) with a G2V spectral type for the stars, similarly to Hoyer et al. (2023), and we set the position at the center of the CCD, where the nucleus was observed. It has been shown, however, that a change of spectral type has a mild impact on the NSC modeling (Böker et al. 2004; Georgiev & Böker 2014).
@@
The errors on the Sérsic parameters were derived for each reliably modeled NSC by injecting artificial NSCs in the HST residual images. The artificial NSCs were modeled using the python wrapper of the software imfit (Erwin 2015) to which we set the parameters of the Sérsic model to be that of the considered NSC. Each artificial source was injected at a random position on the CCD where the target galaxy was centered. In the case of a double nucleated galaxy, we computed the error of each NSC separately. To avoid superposition with bright sources, we isolated the injected NSC from these based on a segmentation image from source extractor. We applied the same fitting method to the artificial sources as for the NSCs. To estimate the error on a given parameter, we ran galfit on 100 artificial NSCs leaving only this parameter free to change, and computed the standard deviation of the 100 resulting models.
@@
We obtained a reliable Sérsic model for 14 and three NSCs in the F606W and F814W filter, respectively, and for 14 NSCs in both filters. When a model is available for both F606W and F814W images, the parameters tend to agree within the computed errors. Therefore, when applicable, we use the average value of the derived Re,NSC, , and ellipticity. All the derived properties and their uncertainties are reported in Table 1. Considering the spatial resolution of the observations and the range of distance covered by the host galaxies, the NSCs in the nearest[farthest] galaxies should have R pc to start being resolved, increasing the difficulty to marginally resolve the NSCs in the farthest galaxies. Considering galaxies farther than 20 Mpc, 16 NSCs are close to be unresolved and have effective radius R pix. Those NSCs are highlighted in Table 1. In addition, we were unable to derive the structural parameters of the NSC of MATLAS-1630, MATLAS-262, and MATLAS-627, all located at a distance between 30 Mpc and 46 Mpc. The remaining galaxies without a NSC model show an intricate central region, consisting of either a possible NSC with stellar tails, or multiple bright clusters close the center. The dwarfs with such complex NSCs, including the double nucleated dwarfs, are studied in detail in Poulain et al. (2025).
@@ 5.2 Color and stellar mass
We derived the color of the NSCs from the total magnitudes of the Sérsic modeling when a reliable model was available in both bands. Otherwise, in case of a one-band-only reliable model, we estimated the color based on aperture magnitude, with a varying aperture size proportional to Re,NSC. We measured the aperture magnitudes making use of the python package photutils (Bradley et al. 2020), and applied the aperture corrections reported for the ACS/WFC in Bohlin (2016). Comparing the obtained colors from aperture and model measurements, we found the best agreement for an aperture radius of 2Re,NSC, with a median difference between the two computed colors of 0.04 and a maximum difference below 0.1. For the 12 NSCs without a reliable model, we used the derived photometry from either Poulain et al. (2025) for the complex NSCs, or Marleau et al. (2024) for the others, when available.
We converted the magnitudes to the system, applying the same approach as Hoyer et al. (2023), and employed the synthetic color transformation coefficients of Sirianni et al. (2005). In the case of a NSC with only one magnitude available, for example , we defined as with being the color derived from aperture photometry. We corrected the obtained and -band magnitudes for Galactic extinction using the estimates from Schlafly & Finkbeiner (2011) at the position of the NSCs.
We computed the stellar mass of the NSCs from the relation between the stellar mass-to-light ratio in the band and the color defined based on the work and data from Roediger & Courteau (2015):
For consistency, we also derived the stellar mass of the dwarfs according to the work of Roediger & Courteau (2015) using the structural properties and photometry of the dwarfs measured on the MATLAS images published in Poulain et al. (2021) and Poulain et al. (2022), making use of the stellar mass-to-light ratio in the band and the color relation:
We find the obtained stellar masses to be on average 0.35 dex less massive than those from Poulain et al. (2021). This difference is consistent with the findings of Habas et al. (2020), and is due to the use of Bell et al. (2003) method to derive the stellar masses in Poulain et al. (2021), which seems to generate the largest mass difference as compared to other methods (Roediger & Courteau 2015). Thus, in Section 6.3, we will correct the stellar masses derived with Bell et al. (2003) when comparing our galaxies to other samples.
To estimate the uncertainties on the stellar masses, we propagated the errors on the photometry from the artificial sources for the Sérsic models and from photutils for aperture magnitude measurements. We note that, in addition to these errors, scatters in the relations between the mass-to-light ratios and colors will add an average of 0.3 dex uncertainty on the estimated stellar mass (Roediger & Courteau 2015).
@@ 6 Structural and photometric properties
In this section, we present the structural properties of the 31 NSCs with a reliable Sérsic model, as well as the photometry of 41 NSCs. We first compare them to NSC properties from the literature, and then detail their color profile.
@@ 6.1 Catalogs of NSCs
For comparison purposes, we compiled the results of several NSC studies targeting both dwarfs and more massive galaxies located in various environments. We used two types of samples: a structural parameters-based and a photometry-based sample. The structural parameters-based sample gathers the properties of NSCs observed by HST and is used to study their Re,NSC, ellipticity and Sérsic index as a function of the NSCs absolute magnitude and stellar mass. In the LV, Pechetti et al. (2020)333We only considered the best quality NSC models (0 and 1) to avoid galaxies highly contaminated by dust. and Hoyer et al. (2023) recently presented Sérsic modeling of NSCs in galaxies with stellar masses between and M⊙. In the dense environment of galaxy clusters, a first analysis of the size and luminosity of NSCs in elliptical galaxies was performed based on the HST data from the ACS Fornax and Virgo cluster surveys (Turner et al. 2012; Côté et al. 2006). We also compare our NSCs to those located in field LTGs at distances up to 40 Mpc (Georgiev & Böker 2014; Georgiev et al. 2016).
The type of model fitted to the underlying galaxy light profile impacts the reliability of the derived NSC properties, especially in massive galaxies, as a single Sérsic profile is sometimes used instead of a multi-component surface brightness profile in the presence of a bulge, bar, or disk (Peng et al. 2002). Because of the use of a single Sérsic or a core-Sérsic (Graham & Guzmán 2003) profile to model the nucleated galaxies, we cleaned-up the samples of the ACS Fornax and Virgo cluster surveys from objects that would have required a multi-component fitting. In the case of Côté et al. (2006), we removed the galaxies with disks or bars identified in Ferrarese et al. (2006), while we removed the galaxies from Turner et al. (2012) brighter than a total blue magnitude of 13.5 mag or described as better fitted by a multi-Sérsic profile.
The photometry-based sample is composed of the measured photometry and stellar masses of the NSCs and their host galaxy. It combines the NSCs from the structural parameters-based sample to ground-based observations. We added the NSC population of dwarf galaxies surrounding LTGs from Carlsten et al. (2022), as well as the photometric measurements and derived stellar masses from Sánchez-Janssen et al. (2019) and Spengler et al. (2017) in the Virgo cluster, and from Ordenes-Briceño et al. (2018) and Su et al. (2022) in Fornax.
@@ 6.2 Relations between structural properties
In Figure 3, we present several relations between the structural properties of the NSCs. We first investigate the variation of the Re,NSC with MV,NSC and the stellar mass . Considering here only studies where NSCs were modeled, we compare the MATLAS NSCs to previously presented samples. When the stellar mass was not available, we used the available photometry and mass-to-light ratios from Roediger & Courteau (2015) to calculate it ourselves. Several studies showed that NSCs with M⊙ or Mmag tend to be larger as they get brighter and more massive (e.g., Georgiev & Böker 2014; Norris et al. 2014; Neumayer et al. 2020; Hoyer et al. 2023), and that this trend appears to break for fainter, less massive NSCs. We show the relation fitted by Georgiev & Böker (2014) in the top left panel and see that, while the brightest MATLAS NSCs with Mmag seem to follow the relation, we do not see an effect of the luminosity or mass on Re,NSC when considering the whole sample. We show in the top right panel of Figure 3 the relation R M found for star clusters formed from GC mergers in the simulations from Bekki et al. (2004). Turner et al. (2012) pointed out that the relation should evolve depending on the number of mergers that the NSC undergoes, with the most massive ones following a steeper relation such that R M∗, which we also added to the plot. While some NSCs from our sample might follow either of these two relations, we cannot confirm the assumption that these NSCs formed from the migration scenario based on their Re,NSC, mass, and luminosity only.
@@
In the bottom left panel, we investigate a potential relation between the ellipticity and the stellar mass of the NSCs. Spengler et al. (2017) and Hoyer et al. (2023) found that the brighter, and massive the NSCs, respectively, the flatter they are. However, we do not see such a trend for the MATLAS dwarfs, LTGs and combined LV samples. On the contrary, the bright MATLAS NSCs tend to be rounder than the faint ones. But, given the large errors on the ellipticity, we cannot affirm any relation.
@@
Finally, we show the Sérsic index as a function of the stellar mass. In the LV, a weak anti-correlation was reported and fitted by Pechetti et al. (2020), in agreement with the structural properties found for bright NSCs in nearby LTGs by Carson et al. (2015). Adding data for NSCs in dwarf galaxies, Hoyer et al. (2023) performed another regression resulting in a less steep slope. This difference of relation might however be driven by the high values, as Pechetti et al. (2020) considered only NSCs with , while Hoyer et al. (2023) used the whole sample. We indicate this cut in the plot. Unlike the findings in the LV, the massive MATLAS NSCs seem to have a larger Sérsic index than the low-mass ones. This correlation is consistent with the observations of Spengler et al. (2017) for NSCs in the Virgo cluster. Yet, we can see that the MATLAS NSCs fit into the large scatter found for the LV population. Thus, considering both the LV and MATLAS samples, it is hard to find any relation. In terms of stellar distributions, a higher Sérsic index means a population that gets more concentrated towards the center of the NSC within Re,NSC when : Carson et al. (2015) reported that Sérsic profiles with indicate a difference on the extent of the profile beyond Re,NSC, rather than a difference in central stellar concentration.
6.3 NSCs-to-galaxy stellar mass relation

### v2

@@ 3 HST observations
For this study a subsample of 79 of the 2210 dwarf galaxies identified in the MATLAS survey were observed with the HST Advanced Camera for Surveys (ACS) in the Wide Field Channel (WFC) mode. The HST observations were performed during the Cycle 27 program GO-16082 (PI: O. Müller), as well as the Cycle 28 and 29 snapshot programs SNAP-16257 and SNAP-16711 (PI: F. Marleau). We used the reduced charge transfer efficiency corrected (CTE-corrected) drc drizzled images produced by the STScI standard pipeline, which have a pixel size of 0.05 arcsec/pixel. Each dwarf is centered in one of the charge coupled devices (CCDs) to optimize the background estimate and ensure that no object is lost (e.g., GCs) in the chip gap or outside the field of view. The galaxies were observed in the F606W and F814W filters with an exposure time in each filter of 1030 s for MATLAS-2019, and 824 s for the others. We computed the VEGAmag zero-points from the python module acszpt222https://acstools.readthedocs.io/en/latest/acszpt.html made by the ACS Team. The targets of the follow-up observations were selected to study the GC population of the MATLAS UDGs (e.g., Mueller2020; Mueller2021; Marleau2021; Marleau2024; Pfeffer2024; Buzzo2025; Haacke2025). Thus, our sample is mainly composed of low surface brightness galaxies and large dwarf galaxies (see Fig. 2 in Marleau2024).
@@ 4 The nucleated dwarf sample
In Poulain2021, we defined a nucleus as a compact source located up to 0.5 effective radius from the photocenter of the dwarf, and that is the brightest compact source within one effective radius. Based on this definition, and after removing one likely foreground star contaminant using the Gaia DR2 catalog (GaiaDR2), 507 nucleated dwarfs were identified in the MATLAS dwarf galaxy sample. We note, however, that of the nucleus sample are located within a galactocentric distance of 0.1 effective radius, or 0.83″, a value lower than the observed PSF of the ground-based observation of the MATLAS data. For 16 nucleated dwarfs, up to three compact sources of similar magnitude were located close to the photocenter and were brighter than the surrounding point sources. In these cases, we defined those galaxies as multinucleated.
Of the 507 identified MATLAS nucleated dwarfs, 30 have HST observations. However, because the ACS camera has a significantly higher resolution than the MegaCam camera, we were able to marginally resolve the nuclei, leading to the rejection of two nuclei that were actually background sources, and the identification of 15 new nuclei. Two of these are newly resolved double nuclei located in already flagged nucleated galaxies. This results in a sample of 41 nucleated dwarfs with HST data to be studied. In Fig. 1 we show a comparison between the resolution of the central region from the HST and MATLAS images for 6 nucleated dwarfs, including three of these new nuclei; color cutouts of the 41 nucleated dwarfs observed with HST are displayed in Fig. 2. In terms of morphology, 40 of the nucleated dwarfs were classified as dE during the initial classification (Habas2020), and only one, MATLAS-682, was defined as a dI.
@@ 5.1 Sérsic modeling
In order to obtain the structural parameters and photometry of the NSCs, we divided the modeling process into two parts. The aim of the first step was to remove the smooth light profile of the dwarf galaxy from the image. As the background is not constant all over the HST image, we estimated the local background of the galaxy. We computed the median sky from clipping in an annulus of radius 15″ on the outskirts of the galaxy where the light profile stops decreasing to become constant. To optimize the galaxy modeling, we masked all nuclei with a circle of radius 1″, a good compromise to cover each NSC without masking a too large region of the galaxies. If necessary, we manually created an additional mask of surrounding bright sources, such as galaxies and stars, with the help of saoimageds9 (DS92000). Finally, we ran galfit (Peng2010) to apply a Sérsic profile (Sersic1963) to the dwarf galaxy, fixing the value of the sky to the estimate. The Sérsic model has long been used in the literature for the NSC modeling (e.g., Graham2009; Seth2010; Turner2012; Lyubenova2013; Carson2015; Spengler2017). Because of the exposure time of the observations, it can be very difficult to differentiate the observed galaxies from the background given their low surface brightness. As a consequence, the Sérsic models systematically have a smaller Re than the models on the CFHT images (Poulain2021), as the flux in the outer part of the galaxies blends into the sky noise. We also obtained less accurate photometric measurements, especially for the faintest galaxies. As a consequence, when the galaxies structural properties and photometry were studied, we used the values from Poulain2021 based on the CFHT images. The second step focused on the Sérsic modeling of the NSCs. For this, we created a cutout of size 5″ centered on the nucleus from the galaxy cleaned image, i.e., the residual image from galfit. The PSF varies with the position on the CCD and the stellar population of the stars; thus, to precisely compute the structural properties of the NSCs, we generated the PSF of each NSC from tinytim (Kirst2011) with a G2V spectral type for the stars, similarly to Hoyer2023, and we set the position at the center of the CCD, where the nucleus was observed. It has been shown, however, that a change of spectral type has a mild impact on the NSC modeling (Boeker2004; Georgiev2014).
@@
The errors on the Sérsic parameters were derived for each reliably modeled NSC by injecting artificial NSCs in the HST residual images. The artificial NSCs were modeled using the python wrapper of the software imfit (Erwin2015) to which we set the parameters of the Sérsic model to be that of the considered NSC. Each artificial source was injected at a random position on the CCD where the target galaxy was centered. In the case of a double nucleated galaxy, we computed the error of each NSC separately. To avoid superposition with bright sources, we isolated the injected NSC from these sources based on a segmentation image from source extractor. We applied the same fitting method to the artificial sources as for the NSCs. To estimate the error on a given parameter, we ran galfit on 100 artificial NSCs leaving only this parameter free to change, and computed the standard deviation of the 100 resulting models.
@@
We obtained a reliable Sérsic model for 14 and 3 NSCs in the F606W and F814W filter, respectively, and for 14 NSCs in both filters. When a model is available for both F606W and F814W images, the parameters tend to agree within the computed errors. Therefore, when applicable, we used the average value of the derived Re,NSC, , and ellipticity. All the derived properties and their uncertainties are reported in Table 1. Considering the spatial resolution of the observations and the range of distances covered by the host galaxies, the NSCs in the nearest (resp. farthest) galaxies should have R pc (resp. 10 pc) to start being resolved, increasing the difficulty to marginally resolve the NSCs in the farthest galaxies. Considering galaxies farther than 20 Mpc, 16 NSCs are close to unresolved and have effective radius R pix. Those NSCs are highlighted in Table 1. In addition, we were unable to derive the structural parameters of the NSC of MATLAS-1630, MATLAS-262, and MATLAS-627, all located at a distance between 30 Mpc and 46 Mpc. The remaining galaxies without a NSC model show an intricate central region, consisting of either a possible NSC with stellar tails or multiple bright clusters close the center. The dwarfs with such complex NSCs, including the double nucleated dwarfs, are studied in detail in Poulain2025.
@@ 5.2 Color and stellar mass
We derived the color of the NSCs from the total magnitudes of the Sérsic modeling when a reliable model was available in both bands. Otherwise, in the case of a one-band-only reliable model, we estimated the color based on aperture magnitude, with a varying aperture size proportional to Re,NSC. We measured the aperture magnitudes making use of the python package photutils (Bradley2020), and applied the aperture corrections reported for the ACS/WFC in Bohlin2016. Comparing the obtained colors from aperture and model measurements, we found the best agreement for an aperture radius of 2Re,NSC, with a median difference between the two computed colors of 0.04 and a maximum difference below 0.1. For the 12 NSCs without a reliable model, we used the derived photometry from either Poulain2025 for the complex NSCs or Marleau2024 for the others, when available.
We converted the magnitudes to the system, applying the same approach as Hoyer2023, and employed the synthetic color transformation coefficients of Sirianni2005. In the case of a NSC with only one magnitude available, for example , we defined as , with being the color derived from aperture photometry. We corrected the obtained - and -band magnitudes for Galactic extinction using the estimates from Schlafly2011 at the position of the NSCs.
We computed the stellar mass of the NSCs from the relation between the stellar mass-to-light ratio in the -band and the color defined based on the work and data from Roediger2015:
For consistency, we also derived the stellar mass of the dwarfs according to the work of Roediger2015 using the structural properties and photometry of the dwarfs measured on the MATLAS images published in Poulain2021 and Poulain2022, making use of the stellar mass-to-light ratio in the -band and the color relation:
On average, we find the obtained stellar masses to be 0.35 dex less massive than those from Poulain2021. This difference is consistent with the findings of Habas2020, and is due to the use of the Bell2003 method to derive the stellar masses in Poulain2021, which seems to generate the largest mass differences compared to other methods (Roediger2015). Thus, in Sect. 6.3 we correct the stellar masses derived with Bell2003 when comparing our galaxies to other samples.
To estimate the uncertainties on the stellar masses, we propagated the errors on the photometry from the artificial sources for the Sérsic models and from photutils for aperture magnitude measurements. We note that, in addition to these errors, scatters in the relations between the mass-to-light ratios and colors will add an average of 0.3 dex uncertainty on the estimated stellar mass (Roediger2015).
@@ 6 Structural and photometric properties
In this section we present the structural properties of the 31 NSCs with a reliable Sérsic model, as well as the photometry of 41 NSCs. We first compare them to NSC properties from the literature, and then detail their color profile.
@@ 6.1 Catalogs of NSCs
For comparison purposes, we compiled the results of several NSC studies targeting both dwarfs and more massive galaxies located in various environments. We used two types of samples: structural parameter-based and photometry-based. The structural parameter-based sample gathers the properties of NSCs observed by HST and is used to study their Re,NSC, ellipticity, and Sérsic index as a function of the NSCs absolute magnitude and stellar mass. In the LV, Pechetti2020333We only considered the best-quality NSC models (0 and 1) to avoid galaxies highly contaminated by dust. and Hoyer2023 recently presented Sérsic modeling of NSCs in galaxies with stellar masses between and M⊙. In the dense environment of galaxy clusters, a first analysis of the size and luminosity of NSCs in elliptical galaxies was performed based on the HST data from the ACS Fornax and Virgo cluster surveys (Turner2012; Cote2006). We also compare our NSCs to those located in field LTGs at distances up to 40 Mpc (Georgiev2014; Georgiev2016).
The type of model fitted to the underlying galaxy light profile impacts the reliability of the derived NSC properties, especially in massive galaxies, as a single Sérsic profile is sometimes used instead of a multicomponent surface brightness profile in the presence of a bulge, bar, or disk (Peng2002). Because of the use of a single-Sérsic or a core-Sérsic (Graham2003) profile to model the nucleated galaxies, we cleaned up the samples of the ACS Fornax and Virgo cluster surveys from objects that would have required a multicomponent fitting. In the case of Cote2006, we removed the galaxies with disks or bars identified in Ferrarese2006, while we removed the galaxies from Turner2012 brighter than a total blue magnitude of 13.5 mag or described as better fitted by a multi-Sérsic profile.
The photometry-based sample is composed of the measured photometry and stellar masses of the NSCs and their host galaxy. It combines the NSCs from the structural parameter-based sample to ground-based observations. We added the NSC population of dwarf galaxies surrounding LTGs from Carlsten2022, as well as the photometric measurements and derived stellar masses from Janssen2019 and Spengler2017 in the Virgo cluster, and from Ordenes2018 and Su2022 in Fornax.
@@ 6.2 Relations between structural properties
In Fig. 3 we present several relations between the structural properties of the NSCs. We first investigated the variation of the Re,NSC with MV,NSC and the stellar mass . Considering only studies where NSCs were modeled, we compared the MATLAS NSCs to previously presented samples. When the stellar mass was not available, we used the available photometry and mass-to-light ratios from Roediger2015 to calculate it ourselves. Several studies showed that NSCs with M⊙ or Mmag tend to be larger as they get brighter and more massive (e.g., Georgiev2014; Norris2014; Neumayer2020; Hoyer2023), and that this trend appears to break for fainter, less massive NSCs. We show the relation fitted by Georgiev2014 in the top left panel and see that, while the brightest MATLAS NSCs with Mmag seem to follow the relation, we do not see an effect of the luminosity or mass on Re,NSC when considering the whole sample. We show in the top right panel of Fig. 3 the relation R M found for star clusters formed from GC mergers in the simulations from Bekki2004. Turner2012 pointed out that the relation should evolve depending on the number of mergers that the NSC undergoes, with the most massive ones following a steeper relation such that R M∗, which we also added to the plot. While some NSCs from our sample might follow either of these two relations, we cannot confirm the assumption that these NSCs formed from the migration scenario based on their Re,NSC, mass, and luminosity only.
@@
In the bottom left panel we investigate a potential relation between the ellipticity and the stellar mass of the NSCs. Spengler2017 and Hoyer2023 found that the brighter and more massive the NSCs, respectively, the flatter they are. However, we do not see such a trend for the MATLAS dwarfs, LTGs, and combined LV samples. On the contrary, the bright MATLAS NSCs tend to be rounder than the faint ones, but given the large errors on the ellipticity, we cannot affirm any relation.
@@
Finally, we show the Sérsic index as a function of the stellar mass. In the LV, a weak anti-correlation was reported and fitted by Pechetti2020, in agreement with the structural properties found for bright NSCs in nearby LTGs by Carson2015. Adding data for NSCs in dwarf galaxies, Hoyer2023 performed another regression resulting in a less steep slope. However, this difference of relation might be driven by the high values as Pechetti2020 considered only NSCs with , while Hoyer2023 used the whole sample. We indicate this cut in the plot. Unlike the findings in the LV, the massive MATLAS NSCs seem to have a larger Sérsic index than the low-mass ones. This correlation is consistent with the observations of Spengler2017 for NSCs in the Virgo cluster. However, we can see that the MATLAS NSCs fit into the large scatter found for the LV population. Thus, considering both the LV and MATLAS samples, it is hard to find any relation. In terms of stellar distributions, a higher Sérsic index means a population that becomes more concentrated toward the center of the NSC within Re,NSC when ; Carson2015 reported that Sérsic profiles with indicate a difference in the extent of the profile beyond Re,NSC, rather than a difference in central stellar concentration.
6.3 NSC-to-galaxy stellar mass relation